OpenStudy (anonymous):

could someone please help with 4 of my questions? im not sure what im doing incorrectly! :(

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −x − 2 y = 2x − 2 y = −x − 7 2. Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −x + 3 y = −x + 6 y = 2x + 3 3. Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −x − 6 y = −x − 14 y = 3x − 6 y = 3x − 14 4. Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = x − 10 y = x + 10 y = −x − 10 y = −x + 10

jimthompson5910 (jim_thompson5910):

please ask one at a time

jimthompson5910 (jim_thompson5910):

I'll help with the first one

jimthompson5910 (jim_thompson5910):

y = 2x + 2 has a slope of 2 anything parallel to this will also have a slope of 2

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so m = 2 y = mx + b y = 2x + b .. plug in the slope m = 2 3 = 2(5) + b ... plug in the point the line goes through, (5,3) 3 = 10 + b 3 - 10 = b -7 = b b = - 7 so we know m = 2 and b = -7

jimthompson5910 (jim_thompson5910):

which means the parallel line is y = 2x - 7

OpenStudy (anonymous):

thank you for the first one :) i see what i did wrong on that one

jimthompson5910 (jim_thompson5910):

hopefully that will help you with the rest if not, then let me know

OpenStudy (anonymous):

im still having a bit of trouble :/ sorry

jimthompson5910 (jim_thompson5910):

its ok

jimthompson5910 (jim_thompson5910):

the slope of y = 2x + 2 is m = 2 flip the fraction and flip the sign to get m = -1/2

jimthompson5910 (jim_thompson5910):

so the slope of the perpendicular line is m = -1/2

OpenStudy (anonymous):

ok got that part

jimthompson5910 (jim_thompson5910):

y = mx + b y = (-1/2)x + b 3 = (-1/2)*6 + b solve for b

jimthompson5910 (jim_thompson5910):

what do you get when you do?

OpenStudy (anonymous):

got -36?

jimthompson5910 (jim_thompson5910):

hmm a bit off

jimthompson5910 (jim_thompson5910):

3 = (-1/2)*6 + b 3 = -6/2 + b 3 = -3 + b b = ??

OpenStudy (anonymous):

-1?

jimthompson5910 (jim_thompson5910):

no you have to add 3 to both sides now

OpenStudy (anonymous):

lol sorry... 6?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the equation is y = -1/2x + 6

OpenStudy (anonymous):

ughh... i totally dont get these... there so hard for me.. idk why!

jimthompson5910 (jim_thompson5910):

keep at it and hopefully things will click better

OpenStudy (anonymous):

only have 3 more then im down... ; p

jimthompson5910 (jim_thompson5910):

try the next one and tell me what you get